Question on Profit & Loss:
 
A Florist bought some roses.He paid one Rs for 7 roses. He repacked the roses in bouquet of 5 each. Each bouquet was sold for 1 Rs and he sold all flowers. Calculate his profit % or Loss %

Grade:

1 Answers

Nishant Vora IIT Patna
9 years ago
Hi student,

Solution is easy.
Cost price is 1rs/ 7roses
So cost price of one rose is 1/7rs

Now sell price is 1rs/5roses
So sell price of 1 rose= 1/5 rs

Profit = SP – CP
So, Profit = 1/5 – 1/7 = 2/35rs ( this is profit in rs)

CP Profit
1/7 2/35
100 ?

So by unitary method
Profit(%) = 200*7/35 = 40%